For a school's computer network, the event of a network failure can be modeled as a Poisson process at a rate of .1 failures/day.

Find the probability that the next network failure will occur within 5 days.
Answer #1

As the event of a network failure is poisson distributed, the interarrival time is exponential distributed.

Probability that the next network failure will occur within 5 days = P(T <= 5)

The formula for the same is:

λ = 0.1 = events/day

P(T<=5) = 1 - e-0.1*5 = 0.394